# A Study of Fairness Concerns in AI-based Mobile App Reviews

### v2

@@ 9. Conclusion and Future Work
In this study, we analyzed the reviews of AI-based apps to identify the types of fairness concerns users experience when using AI-based apps. We also investigated root causes reported by app owners to justify the fairness concerns. Our analysis was based on $\approx$9.5M reviews collected from 108 AI-based apps. Given that app reviews cover different topics, we first developed a set of ML and DL binary classifiers to identify reviews that include discussions around fairness (i.e., fairness reviews). We used a manually constructed dataset of fairness and non-fairness reviews to train and evaluate these ML and DL classifiers. We found that Logistic Regression with TF-IDF, USE, and Word2Vec as the feature extraction technique reaches the highest precision of 0.94. Furthermore, our experiments with K-mean clustering on $\approx$92K fairness reviews resulted in six types of fairness concerns in AI-based apps. Our qualitative analysis of 2,248 app owners’ responses identified six root causes (e.g., ‘policy issues’, ‘development cost’) of the fairness concerns.
We plan to conduct further studies with other stakeholders of AI-based apps such as app owners, developers, and policymakers, and obtain more data from other sources (e.g., GitHub) to generalize our understanding of fairness concerns in AI-based apps. We also plan to experiment with other clustering techniques.

### v3

@@ 9. Conclusion and Future Work
In this study, we analyzed the reviews of AI-based apps to identify the types of fairness concerns users experience when using AI-based apps. We also investigated root causes reported by app owners to justify the fairness concerns. Our analysis was based on $\approx$9.5M reviews collected from 108 AI-based apps. Given that app reviews cover different topics, we first developed a set of ML and DL models to identify reviews that include discussions around fairness (i.e., fairness reviews). We used a manually constructed dataset of fairness and non-fairness reviews to train and evaluate these ML and DL models. We found that Logistic Regression with TF-IDF, USE, and Word2Vec as the feature extraction technique reaches the highest precision of 0.94. Furthermore, our experiments with K-mean clustering on $\approx$92K fairness reviews resulted in six types of fairness concerns in AI-based apps. Our qualitative analysis of 2,248 app owners’ responses identified six root causes (e.g., ‘policy issues’, ‘development cost’) of the fairness concerns.
We plan to conduct further studies with other stakeholders of AI-based apps such as app owners, developers, and policymakers, and obtain more data from other sources (e.g., GitHub) to generalize our understanding of fairness concerns in AI-based apps. We also plan to experiment with other clustering techniques and evaluate the extent to which the employed learning models adhere to ethical principles, with a particular focus on fairness.